how do i verify and prove that
sin(beta - 3[pi]/2) = cos (beta)

Answer by lwsshak3(6518)

how do i verify and prove that
sin(beta - 3[pi]/2) = cos (beta)
**
Use x in place of beta
starting with left side
use addition formula for sin
sin(x-3π/2)=[sinx*cos(3π/2)]-[cosx*sin(3π/2)]
cos(3π/2)=0
sin(3π/2)=-1
sin(x-3π/2)=[sinx*0]-[cosx*(-1)]=0-(-cosx)=cosx
Verified:
left side=right side